### v1

@@ 7 Model-dependence of DM-galaxy offsets
In this section we discuss implications of the results obtained from the simulation of galaxy cluster mergers in the previous section. In particular, a main quantity of interest in this context is the offset between the DM halo and either the galaxy population or the brightest galaxy of the cluster (i.e. the BCG). Here we discuss our findings on how these offsets depend on the underlying model describing DM self-interactions. Specifically, our focus is on the characteristic angle-dependence of the differential cross section as predicted in realistic (e.g. light mediator) models, that we take into account in the simulations based on the hSIDM scheme developed in this work.
@@ 7.1 Dependence on viscosity cross section for fixed anisotropy
Figure 12: Dependence of the DM peak positions (left) as well as maximal DM-galaxy (upper right) and DM-BCG (lower right) offsets on the viscosity cross section $\sigma_{\mathrm{V}}/m_{\chi}$, while keeping the amount of anisotropy of the differential cross section $\mathrm{d}\sigma/\mathrm{d}\Omega$ fixed (‘Møller’ cross section Eq. (1) with $r=10^{4}$). The time-axis on the left is rescaled such as to align the first and second pericentre passage for all models. The DM peaks remain the closer to each other the larger $\sigma_{\mathrm{V}}/m_{\chi}$. The offsets increase (roughly linearly) with $\sigma_{\mathrm{V}}/m_{\chi}$ within the considered range (right column). We note that our results are independent of the choice of critical angle $\theta_{\mathrm{c}}$ entering the hSIDM scheme, with offsets for $\theta_{\mathrm{c}}=0.1$ and $0.3$ being compatible with each other. The error bars are computed following Fischer et al. (2021b).
As a first step, we verify the expectation that the offsets increase when increasing the overall DM self-interaction rate. To demonstrate this dependence, we simulate galaxy cluster mergers for models with various values of the viscosity cross section in the range $0.2\,{\rm cm}^{2}\,{\rm g}^{-1}\leq\sigma_{\mathrm{V}}/m_{\chi}\leq 3\,{\rm cm}^{2}\,{\rm g}^{-1}$, but keeping the angle-dependence of the cross section fixed (we choose a ‘Møller’ cross section Eq. (1) with $r=10^{4}$). The time-evolution of the DM peak positions for the two halos is shown in Fig. 12 (left panel). Compared to collisionless DM (CDM, black dashed lines), the peaks remain the more close to each other the larger $\sigma_{\mathrm{V}}$, as expected. We note that, following Fischer et al. (2021b), we use a linearly re-scaled time variable, $t_{\rm in}$, here, defined such that the first (second) pericentre occurs at $t_{\rm in}=0\ (1)$, respectively. The maximal absolute DM-galaxy and DM-BCG offsets between the first and second pericentre are shown in the right panels of Fig. 12, versus $\sigma_{\mathrm{V}}/m_{\chi}$. We observe that the offsets increase roughly linearly within the range of shown viscosity cross sections. To illustrate the magnitude, we extract an approximate estimate for both DM-galaxy and DM-BCG offsets given by
@@ 7.2 Dependence on anisotropy for fixed viscosity cross section
Figure 13: Dependence of the DM-galaxy (upper row) and DM-BCG (lower row) offsets on the amount of anisotropy of the differential cross section $\mathrm{d}\sigma/\mathrm{d}\Omega$, while keeping the viscosity cross section $\sigma_{\mathrm{V}}/m_{\chi}=1\,{\rm cm}^{2}\,{\rm g}^{-1}$ fixed. We show results obtained within the hSIDM scheme, assuming the ‘Møller’ cross section Eq. (1) with various values of the anisotropy parameter $r$ (see Eq. (3)) related to the ratio of mediator and DM particle masses. For comparison, we also show the cases of purely forward-dominated scattering (fSIDM) and isotropic scattering (isotropic rSIDM), respectively. The left column shows the time-evolution, the the right column the maximal absolute offsets within $t\in[1.5\,\mathrm{Gyr},3\,\mathrm{Gyr}]$. The maximal offsets for models with realistic angle-dependence approach those for purely forward-dominated scattering for $r\gtrsim 10^{2}$ (corresponding to $m_{\phi}/m_{\chi}\lesssim 0.1v/c$, where $v$ is the typical velocity scale), and those for isotropic scattering for $r\lesssim 1$ ($m_{\phi}/m_{\chi}\gtrsim v/c$). The error bars and error bands are computed following Fischer et al. (2021b).
@@
The left panels of Fig. 13 show the time evolution of the DM-galaxy (top) and DM-BCG (bottom) offset, respectively. As expected from Sect. 6, the offsets are negative shortly after the first pericentre passage, since the DM halos were slowed down when passing through each other, while the offsets are positive around the first apocentre, since the galaxy populations are subsequently slowed down by the gravitational force exerted by the DM mass distribution. The overall size of the offsets tend to be the larger the more anisotropic the scattering is, as can be seen by comparing the lines showing the evolution for $r=1,10,10^{2}$, respectively. This is in line with the findings of Fischer et al. (2021a), where offsets for fSIDM were compared to those for isotropic rSIDM. [^18] Note however that the modified transfer cross section was kept fixed in that case, leading somewhat larger differences; compare to Sect. 6.3 and Fig. 11. Indeed, we find that our results for $r=1,10,10^{2}$ yield offsets that typically lie between those for fSIDM and isotropic rSIDM within the numerical accuracy. To highlight this point, we show the maximal absolute offset within the time interval $t\in[1.5\,\mathrm{Gyr},3\,\mathrm{Gyr}]$ in the right panels of Fig. 13, versus the model parameter $r$. These results allow us to infer the relevance of model-dependence, being one of the main results to be discussed next.
@@ 7.3 Discussion
Models of dark matter self-interactions typically feature angle-dependent scattering. Here we use our results to assess in how far the limits of purely forward-dominated (fSIDM) or isotropic (isotropic rSIDM) scattering yield an accurate approximation to a true angle-dependence described by the ‘Møller’ cross section Eq. (1). Its anisotropy is determined by the parameter $r$ (see Eq. (3)), with fSIDM (isotropic rSIDM) being approached in the limiting cases $r\to\infty$ ($r\to 0$), respectively. Here, we discuss for which finite values of $r$ these limits are approached in practice. In particular, we focus on the maximal DM-galaxy and DM-BCG offsets shown in the right column of Fig. 13. We notice several features:
@@
Let us stress that the angle-dependence enters the merger dynamics in general in a complex way, affecting galaxy and DM distributions differently, see Sect. 6.3. Nevertheless, when considering only the DM-galaxy offset, one may ask the following question: How large would the error on the inferred self-interaction cross section be, when assuming an incorrect model for the angle-dependence? Given that the offsets differ by about a factor of two between the limits of small and large $r$, see Fig. 13, and that their dependence on the viscosity cross section is approximately linear, see Sect. 7.1, we estimate that the deviation in $\sigma_{\mathrm{V}}/m_{\chi}$ would amount to about a factor of two as well. For the modified transfer cross section, this number would go up to a factor $\sim 3$, see Sect. 6.3. However, we stress that these estimates should be taken as indicative only. Fischer et al. (2021b) found that the difference in the size of offsets between isotropic and forward-dominated scattering is larger in unequal-mass mergers compared to equal-mass mergers. Accordingly, this would imply a larger estimate for the error than the factor of two for the viscosity cross section in the case of merging galaxy clusters with a significantly unequal mass ratio. Modelling an observed galaxy cluster merger with detailed simulations of various models for dark matter self-interactions and astrophysical settings could reveal more information than when just considering (maximal) offsets based on a peak identification method that is not applicable to observations.
@@ 8 Conclusions
In this work we propose, validate and employ a numerical scheme to efficiently perform $N$-body simulations of SIDM, taking into account the angle-dependence predicted in realistic particle physics models. Most previous works considered either the case of relatively rare scatterings by large angles (‘rSIDM’, typically assuming an isotropic distribution), or very frequent and purely forward-dominated scatterings (‘fSIDM’). The hybrid scheme developed in this work, dubbed ‘hSIDM’, combines the methods underlying both of these approaches, yielding an efficient algorithm for incorporating DM self-interactions described by an (in principle) arbitrary differential cross section $\mathrm{d}\sigma/\mathrm{d}\Omega$.
@@
In a first step, we extensively validate the approach based on an analytically solvable test set-up. In this context, we also derive a quantitative validity criterion for the effective treatment of small-angle scattering by a drag force combined with a diffusive transverse motion. These insights are used to show under which conditions the hSIDM scheme is expected to improve over a naive sampling of scattering angles in each collision of $N$-body particles, and we demonstrate a speed-up by several orders of magnitude in computation time. The theoretically expected speed-up is estimated in terms of a ratio of angle-weighted cross sections in Eq. (16). We check independence of our results of the auxiliary ‘critical angle’ entering the hSIDM scheme, within a validity regime derived from the condition given in Eq. (36).
@@
Next, we apply hSIDM to simulate equal-mass galaxy cluster mergers including DM self-interactions with characteristic angle-dependence predicted by light mediator models. Concretely, we assume a dependence of the scattering probability on the deflection angle given by Eq. (1). In this parameterisation, the model parameter $r$ (being related to the ratio of mediator and dark matter mass, see Eq. (3)) controls the amount of anisotropy, with $r\gg 1$ corresponding to strongly forward-dominated scattering, and $r\ll 1$ to the isotropic limit. We first investigate whether simulations with different $r$ but identical transfer or viscosity cross section can be mapped onto each other, and find that this is in general not the case, especially when jointly considering the time-evolution of the DM distribution, the galaxy population, the position of the BCG, as well as their relative offsets. Nevertheless, the dependence of the offset on $r$ is somewhat weaker when fixing the viscosity cross section as compared to the case when keeping the (modified) transfer cross section fixed.
@@
Finally, we investigate the dependence of the DM-galaxy and DM-BCG offsets on the properties of the model for DM self-interactions. We find that the maximal offset approaches the (larger) value predicted by purely forward-dominated scattering for $r\gtrsim 10^{2}$. This corresponds to a mediator to DM mass ratio $m_{\phi}/m_{\chi}\lesssim 0.1v/c$, where $v$ is the typical velocity scale. Equivalently, this implies $\sigma_{\rm tot}/\sigma_{\mathrm{\widetilde{T}}}>10$ for the ratio of total and modified momentum transfer cross sections. Conversely, the maximal offset converges to the (smaller) value obtained for isotropic self-scattering when $r\lesssim 1$, i.e. $m_{\phi}/m_{\chi}\gtrsim v/c$ or $\sigma_{\rm tot}/\sigma_{\mathrm{\widetilde{T}}}\lesssim 1.1$. Our results thus yield a quantitative criterion for when the fSIDM and rSIDM limits are applicable for extracting offsets. Beyond that, we find that the angle-dependence of DM self-scattering influences the time-evolution of DM and galaxy populations in a distinct way, suggesting that it should be taken into account when modelling observed galaxy cluster mergers with simulations of self-interacting dark matter.
@@
Our work can be extended by considering broader classes of galaxy cluster merger set-ups (unequal mass, non-central) with a more realistic modelling (e.g. including the intra-cluster medium) and taking also the interplay of the angle- and velocity-dependence of the differential cross section into account. Furthermore, the hSIDM approach could be applied to other situations, such as for example the evolution of satellite populations or for studying the approach of gravothermal collapse.
@@ Appendix C Vector calculation for anisotropic scattering
In this rotated frame, the initial velocities in the centre of mass frame are aligned along the $x$-axis. The velocity vectors before (un-primed) and after (primed) scattering read
@@ Appendix E No-recoil case implementation
In the no recoil case the modified drag force [^20] We note that the prefactor differs from Eq. (39) by a factor 4. This arises from the combination of a factor 2 due to no-recoil kinematics and another factor 2 due to the replacement $\sigma_{\mathrm{\widetilde{T}}}\to\sigma_{\mathrm{T}}$ involving a different normalisation factor. is given by
@@ Appendix F Fixed-angle deflection test
Appendix G Multiple scattering and time-step
The sampling of the scattering angle in the rSIDM scheme is based on the assumption that the physical particles represented by the numerical particles do not scatter more than once per pairwise interaction per time-step. This assumption is implicitly made when sampling the scattering angle for the two numerical particles from the differential cross section instead of a distribution that takes multiple scatterings into account such as Eq. (18) for the no-recoil case. In the following, we test which role this assumption plays in the case of our fixed-angle deflection test, see Sect. 5.2.
When considering the pairwise interaction of numerical particles, the size of the time-step affects how many of the represented physical particles would scatter. It also alters how many physical particles scatter multiple times relative to only once or not at all. We note that this refers to (1) scatterings of physical particles with in a single time-step and (2) scatterings of physical particles belonging to the same pair of numerical particles.
To estimate the probabilities for no and one scatter, we assume that the scattering probability is constant over time

### v2

@@ 7 Model dependence of dark matter-galaxy offsets
In this section we discuss implications of the results obtained from the simulation of galaxy cluster mergers in the previous section. In particular, a main quantity of interest in this context is the offset between the DM halo and either the galaxy population or the brightest galaxy of the cluster (i.e. the BCG). Here we discuss our findings on how these offsets depend on the underlying model describing DM self-interactions. Specifically, our focus is on the characteristic angle dependence of the differential cross section as predicted in realistic (e.g. light mediator) models, that we take into account in the simulations based on the hSIDM scheme developed in this work.
@@ 7.1 Dependence on viscosity cross section for fixed anisotropy
Figure 12: Dependence of the DM peak positions (left) as well as maximal DM-galaxy (upper right) and DM-BCG (lower right) offsets on the viscosity cross section $\sigma_{\mathrm{V}}/m_{\chi}$ while keeping the amount of anisotropy of the differential cross section $\mathrm{d}\sigma/\mathrm{d}\Omega$ fixed (‘Møller’ cross section Eq. (1) with $r=10^{4}$). The time-axis on the left is rescaled such as to align the first and second pericentre passage for all models. The DM peaks remain the closer to each other the larger $\sigma_{\mathrm{V}}/m_{\chi}$. The offsets increase (roughly linearly) with $\sigma_{\mathrm{V}}/m_{\chi}$ within the considered range (right column). We note that our results are independent of the choice of critical angle $\theta_{\mathrm{c}}$ entering the hSIDM scheme, with offsets for $\theta_{\mathrm{c}}=0.1$ and $0.3$ being compatible with each other. The error bars are computed following Fischer et al. (2021b).
As a first step, we verify the expectation that the offsets increase when increasing the overall DM self-interaction rate. To demonstrate this dependence, we simulate galaxy cluster mergers for models with various values of the viscosity cross section in the range $0.2\,{\rm cm}^{2}\,{\rm g}^{-1}\leq\sigma_{\mathrm{V}}/m_{\chi}\leq 3\,{\rm cm}^{2}\,{\rm g}^{-1}$, but keeping the angle dependence of the cross section fixed (we choose a ‘Møller’ cross section Eq. (1) with $r=10^{4}$). The time evolution of the DM peak positions for the two halos is shown in Fig. 12 (left panel). Compared to collisionless DM (CDM, black dashed lines), the peaks remain the more close to each other the larger $\sigma_{\mathrm{V}}$, as expected. We note that, following Fischer et al. (2021b), we use a linearly re-scaled time variable, $t_{\rm in}$, here, defined such that the first (second) pericentre occurs at $t_{\rm in}=0\ (1)$, respectively. The maximal absolute DM-galaxy and DM-BCG offsets between the first and second pericentre are shown in the right panels of Fig. 12, versus $\sigma_{\mathrm{V}}/m_{\chi}$. We observe that the offsets increase roughly linearly within the range of shown viscosity cross sections. To illustrate the magnitude, we extract an approximate estimate for both DM-galaxy and DM-BCG offsets given by
@@ 7.2 Dependence on anisotropy for fixed viscosity cross section
Figure 13: Dependence of the DM-galaxy (upper row) and DM-BCG (lower row) offsets on the amount of anisotropy of the differential cross section $\mathrm{d}\sigma/\mathrm{d}\Omega$ while keeping the viscosity cross section $\sigma_{\mathrm{V}}/m_{\chi}=1\,{\rm cm}^{2}\,{\rm g}^{-1}$ fixed. We show results obtained within the hSIDM scheme, assuming the ‘Møller’ cross section Eq. (1) with various values of the anisotropy parameter $r$ (see Eq. (3)) related to the ratio of mediator and DM particle masses. For comparison, we also show the cases of purely forward-dominated scattering (fSIDM) and isotropic scattering (isotropic rSIDM), respectively. The left column shows the time evolution, the right column the maximal absolute offsets within $t\in[1.5\,\mathrm{Gyr},3\,\mathrm{Gyr}]$. The maximal offsets for models with realistic angle dependence approach those for purely forward-dominated scattering for $r\gtrsim 10^{2}$ (corresponding to $m_{\phi}/m_{\chi}\lesssim 0.1v/c$, where $v$ is the typical velocity scale), and those for isotropic scattering for $r\lesssim 1$ ($m_{\phi}/m_{\chi}\gtrsim v/c$). The error bars and error bands are computed following Fischer et al. (2021b).
@@
The left panels of Fig. 13 show the time evolution of the DM-galaxy (top) and DM-BCG (bottom) offset, respectively. As expected from Sect. 6, the offsets are negative shortly after the first pericentre passage, since the DM halos were slowed down when passing through each other, while the offsets are positive around the first apocentre, since the galaxy populations are subsequently slowed down by the gravitational force exerted by the DM mass distribution. The overall size of the offsets tend to be the larger the more anisotropic the scattering is, as can be seen by comparing the lines showing the evolution for $r=1,10,10^{2}$, respectively. This is in line with the findings of Fischer et al. (2021a), where offsets for fSIDM were compared to those for isotropic rSIDM. [^18] We note, however, that the modified transfer cross section was kept fixed in that case, leading somewhat larger differences; compare to Sect. 6.3 and Fig. 11. Indeed, we find that our results for $r=1,10,10^{2}$ yield offsets that typically lie between those for fSIDM and isotropic rSIDM within the numerical accuracy. To highlight this point, we show the maximal absolute offset within the time interval $t\in[1.5\,\mathrm{Gyr},3\,\mathrm{Gyr}]$ in the right panels of Fig. 13, versus the model parameter $r$. These results allow us to infer the relevance of model-dependence, being one of the main results to be discussed next.
@@ 7.3 Discussion
Models of dark matter self-interactions typically feature angle-dependent scattering. Here we use our results to assess in how far the limits of purely forward-dominated (fSIDM) or isotropic (isotropic rSIDM) scattering yield an accurate approximation to a true angle dependence described by the ‘Møller’ cross section Eq. (1). Its anisotropy is determined by the parameter $r$ (see Eq. (3)), with fSIDM (isotropic rSIDM) being approached in the limiting cases $r\to\infty$ ($r\to 0$), respectively. Here, we discuss for which finite values of $r$ these limits are approached in practice. In particular, we focus on the maximal DM-galaxy and DM-BCG offsets shown in the right column of Fig. 13. We notice several features:
@@
Let us stress that the angle dependence enters the merger dynamics in general in a complex way, affecting galaxy and DM distributions differently, see Sect. 6.3. Nevertheless, when considering only the DM-galaxy offset, one may ask how large the error on the inferred self-interaction cross section would be when assuming an incorrect model for the angle dependence. Given that the offsets differ by about a factor of two between the limits of small and large $r$, see Fig. 13, and that their dependence on the viscosity cross section is approximately linear, see Sect. 7.1, we estimate that the deviation in $\sigma_{\mathrm{V}}/m_{\chi}$ would amount to about a factor of two as well. For the modified transfer cross section, this number would go up to a factor of approximately three (see Sect. 6.3). However, we stress that these estimates should be taken as indicative only. Fischer et al. (2021b) found that the difference in the size of offsets between isotropic and forward-dominated scattering is larger in unequal-mass mergers compared to equal-mass mergers. Accordingly, this would imply a larger estimate for the error than the factor of two for the viscosity cross section in the case of merging galaxy clusters with a significantly unequal mass ratio. Modelling an observed galaxy cluster merger with detailed simulations of various models for dark matter self-interactions and astrophysical settings could reveal more information than when just considering (maximal) offsets based on a peak identification method that is not applicable to observations.
@@ 8 Conclusions
In this work, we have proposed, validated, and employed a numerical scheme to efficiently perform $N$-body simulations of SIDM while taking into account the angle dependence predicted in realistic particle physics models. Most previous works considered either the case of relatively rare scatterings by large angles (‘rSIDM’, typically assuming an isotropic distribution) or very frequent and purely forward-dominated scatterings (‘fSIDM’). The hybrid scheme developed in this work, dubbed ‘hSIDM’, combines the methods underlying both of these approaches, yielding an efficient algorithm for incorporating DM self-interactions described by an (in principle) arbitrary differential cross section $\mathrm{d}\sigma/\mathrm{d}\Omega$.
@@
As a first step, we extensively validated the approach based on an analytically solvable test set-up. In this context, we also derived a quantitative validity criterion for the effective treatment of small-angle scattering by a drag force combined with a diffusive transverse motion. The insights were used to show under which conditions the hSIDM scheme is expected to improve over a naive sampling of scattering angles in each collision of $N$-body particles, and we demonstrated a speed-up by several orders of magnitude in computation time. The theoretically expected speed-up was estimated in terms of a ratio of angle-weighted cross sections in Eq. (16). We checked the independence of our results of the auxiliary ‘critical angle’ entering the hSIDM scheme within a validity regime derived from the condition given in Eq. (36).
@@
Next, we applied hSIDM to simulate equal-mass galaxy cluster mergers including DM self-interactions with the characteristic angle dependence predicted by light mediator models. Concretely, we assumed a dependence of the scattering probability on the deflection angle given by Eq. (1). In this parameterisation, the model parameter $r$ (related to the ratio of mediator and DM mass; see Eq. (3)) controls the amount of anisotropy, with $r\gg 1$ corresponding to strongly forward-dominated scattering, and $r\ll 1$ corresponds to the isotropic limit. We first investigated whether simulations with a different $r$ but an identical transfer or viscosity cross section can be mapped onto each other, and we found that this is in general not the case, especially when jointly considering the time evolution of the DM distribution, the galaxy population, the position of the BCG, and their relative offsets. Nevertheless, the dependence of the offset on $r$ is somewhat weaker when fixing the viscosity cross section as compared to the case when keeping the (modified) transfer cross section fixed.
@@
Finally, we investigated the dependence of the DM-galaxy and DM-BCG offsets on the properties of the model for DM self-interactions. We find that the maximal offset approaches the (larger) value predicted by purely forward-dominated scattering for $r\gtrsim 10^{2}$. This corresponds to a mediator to DM mass ratio of $m_{\phi}/m_{\chi}\lesssim 0.1v/c$, where $v$ is the typical velocity scale. Equivalently, this implies $\sigma_{\rm tot}/\sigma_{\mathrm{\widetilde{T}}}>10$ for the ratio of the total and modified momentum transfer cross sections. Conversely, the maximal offset converges to the (smaller) value obtained for isotropic self-scattering when $r\lesssim 1$ (i.e. $m_{\phi}/m_{\chi}\gtrsim v/c$ or $\sigma_{\rm tot}/\sigma_{\mathrm{\widetilde{T}}}\lesssim 1.1$). Our results thus yield a quantitative criterion for when the fSIDM and rSIDM limits are applicable for extracting offsets. Beyond that, we find that the angle dependence of DM self-scattering influences the time evolution of DM and galaxy populations in a distinct way, suggesting that it should be taken into account when modelling observed galaxy cluster mergers with simulations of self-interacting DM.
@@
Our work can be extended by considering broader classes of galaxy cluster merger set-ups (unequal mass, non-central) with a more realistic modelling (e.g. including the intracluster medium) and by also taking the interplay of the angle- and velocity-dependence of the differential cross section into account. Furthermore, the hSIDM approach could be applied to other situations, such as the evolution of satellite populations, for a precision study of the impact of self-interactions with pronounced angle dependence on isolated halos or for studying the approach of gravothermal collapse.
@@ Appendix C Vector calculation for anisotropic scattering
In this rotated frame, the initial velocities in the centre of mass frame are aligned along the $x$-axis. The velocity vectors before (un-primed) and after (primed) scattering read as
@@ Appendix E No-recoil case implementation
In the no recoil case the modified drag force [^20] We note that the prefactor differs from Eq. (39) by a factor of four. This arises from the combination of a factor of two due to no-recoil kinematics and another factor 2 due to the replacement $\sigma_{\mathrm{\widetilde{T}}}\to\sigma_{\mathrm{T}}$ involving a different normalisation factor. is given by
@@ Appendix F Fixed-angle deflection test
Appendix G Multiple scattering and time step
The sampling of the scattering angle in the rSIDM scheme is based on the assumption that the physical particles represented by the numerical particles do not scatter more than once per pairwise interaction per time step. This assumption is implicitly made when sampling the scattering angle for the two numerical particles from the differential cross section instead of a distribution that takes multiple scatterings into account such as Eq. (18) for the no-recoil case. In the following, we test which role this assumption plays in the case of our fixed-angle deflection test, see Sect. 5.2.
When considering the pairwise interaction of numerical particles, the size of the time step affects how many of the represented physical particles would scatter. It also alters how many physical particles scatter multiple times relative to only once or not at all. We note that this refers to (1) scatterings of physical particles with in a single time step and (2) scatterings of physical particles belonging to the same pair of numerical particles.
To estimate the probabilities for no and one scatter, we assume that the scattering probability is constant over time: